The tens digit of a certain two digit number exceeds the units digit by 3. The sum of the digits is 1/7 of the number. Find the number.
Answer by stanbon(75887)

The tens digit of a certain two digit number exceeds the units digit by 3. The sum of the digits is 1/7 of the number. Find the number.
-------------------
Let the number be 10t+u where t is the tens digit and u is the units digit.
EQUATION:
t=u+3
t+u = (1/7)(10t+u)
-------------
Rearrange:
t = u+3
7t + 7u = 10t+u
3t=6u
t = 2u
-----------
2u = u+3
u = 3 (the units digit)
---------
t = u+3 = 3+3 = 6 (the tens digit)
================
Number = 63
=================
Cheers,
San H.


-------------
Substitute to solve for "u":
7(u+3)-6u = 0
u =